Grade upto college level Physical Chemistry

The enthalpy change involved in the oxidation of glucose is 72880 kJ mol-1, Twenty five percent of this energy is available for muscular work. If 100 kJ of muscular work is needed to walk one kilometer, what is the maximum distance that a person will be able to walk after eating 120 g of glucose.

Shane Macguire
12 Years agoGrade upto college level

1 Answer

Deepak Patra
12 Years ago
Hello Student,
Please find the answer to your question
Energy available for muscular work by 1 mole of glucose = 2880 *25/100 = 720 kJ mol-1
Thus 180 g of glucose (mol. Wt. of glucose) supplies 720 kJ
120 g of glucose will supply = 720/180 * 120 = 480 kJ
100 kJ is needed to walk 1 km
Hence, 480 kJ is needed to walk 1/100 * 480 = 4.8 km

Thanks
jitender
askIITians Faculty